A new fluorescent light to replace incandescent bulbs has come on the market. The box says it has an average life of 8,000 hours with a standard deviation of 750 hours. A random sample of 75 light bulbs is taken. Let X be the mean of the sample.
a.) What is the probability that X ≥ 7900?
b.) What is the probability that 7800 ≤ X ≤ 8100?

Answer by stanbon(75887)

A new fluorescent light to replace incandescent bulbs has come on the market. The box says it has an average life of 8,000 hours with a standard deviation of 750 hours. A random sample of 75 light bulbs is taken. Let X be the mean of the sample.
a.) What is the probability that X ≥ 7900?
z(7900) = (7900-8000)/750 = -0.1333
P(x >= 7900) = P(z > -0.1333) = normalcdf(-0.1333,100) = 0.5530
----------------------------------------------------------
b.) What is the probability that 7800 ≤ X ≤ 8100?
Find the z-values of 7800 and of 8100
Then find the probability that z lies between those 2 z-values.
----
Ans: 0.1582
========================
Cheers,
Stan H.